# GraphPilot: Grounded Scene Graph Conditioning for Language-Based Autonomous Driving

### v2

@@ Abstract
Vision-language models have recently emerged as promising planners for autonomous driving, where success hinges on topology-aware reasoning over spatial structure and dynamic interactions from multimodal input. However, existing models are typically trained without supervision that explicitly encodes these relational dependencies, limiting their ability to infer how agents and other traffic entities influence one another from raw sensor data. In this work, we bridge this gap with a novel model-agnostic method that conditions language-based driving models on structured relational context in the form of traffic scene graphs. We serialize scene graphs at various abstraction levels and formats, and incorporate them into the models via structured prompt templates, enabling a systematic analysis of when and how relational supervision is most beneficial. Extensive evaluations on the public LangAuto benchmark show that scene graph conditioning of state-of-the-art approaches yields large and persistent improvement in driving performance. Notably, we observe up to a 15.6% increase in driving score for LMDrive and 17.5% for BEVDriver, indicating that models can better internalize and ground relational priors through scene graph-conditioned training, even without requiring scene graph input at test-time. Code, fine-tuned models, and our scene graph dataset are publicly available at https://github.com/iis-esslingen/GraphPilot.
@@ 1 Introduction
Figure 1: Explicit relational grounding through scene graph conditioning. We visualize four setups, where two binary digits indicate scene graph usage during training and testing: SG00-PT (baseline pretrained, no scene graphs), SG00-FT (baseline fine-tuned without scene graphs), SG10 (scene graphs only during training), and SG11 (scene graphs during training and testing). Models trained under SG10 perform largely on par with SG11, suggesting that they internalize relational structure during training. Performance indicates (mean) driving scores, cf. Tab. 1 and Tab. 4.
@@
Understanding complex traffic scenes remains a central challenge in autonomous driving, especially when decision-making is conditioned on natural language instructions [cui2024survey]. Modern language-based driving agents operate at the intersection of perception, language, and planning, requiring not only accurate scene understanding but also the ability to reason about spatial structure, traffic rules, and interactions among dynamic actors [xie2025vlms-ready]. While recent works have shown that vision-language models (VLMs) can map image and language inputs to plausible trajectories [shao2024lmdrive, winter2025bevdriver, renz2025simlingo, zhang2025vldrive, zhang2025adadrive], their performance in complex environments remains limited. A key limitation is that current models typically rely on representations that do not explicitly encode relational structure, forcing them to infer safety-critical interactions implicitly from dense features [zhang2025safeauto].
@@
Scene graphs provide a structured solution to this limitation [johnson2015image, kassab2024bare]. By explicitly representing a traffic scene as a structured graph of entities and their relations, they encode not only what is present in the environment, but also how entities are related and may influence one another. In scene graphs, nodes represent objects such as vehicles, pedestrians, traffic lights, and roadway elements, while edges express spatial, semantic, and regulatory relations, e.g., which vehicle is in which lane, which traffic light controls which direction, and how nearby actors might interact. This structured relational view enhances semantic understanding and enables more informed planning [chen2025static].
@@
Despite their growing use in indoor navigation and instruction-following tasks [gu2024conceptgraphs, rana2023sayplan, yin2024sgnav, honerkamp2024language, werby2024hierarchical, mohammadi2025more], scene graphs have only recently gained attention in outdoor autonomous driving, mainly for tasks such as risk estimation [malawade2022sg2vec, liu2023learning], and trajectory prediction [jia2023hdgt, zhang2025graphad]. However, their integration into language-based planning remains unexplored. Instead, most existing systems represent the scene using dense geometric or visual features, including image tokens [renz2025simlingo, zhang2025vldrive, zhang2025adadrive] or BEV grids [shao2024lmdrive, winter2025bevdriver, hegde2025dima], followed by generic attention mechanisms. As a result, critical relationships between entities, such as which traffic light controls a given lane or how nearby vehicles might interact, must be inferred implicitly. This gap motivates our central question: Can explicit relational structure of a traffic scene help language-based driving models internalize and reason over spatial and causal dependencies in traffic scenes?
@@
To address this question, we introduce a model-agnostic approach for injecting structured semantic context into language-based autonomous driving systems. We construct traffic scene graphs at each time step and serialize them into human-readable formats (text, JSON, YAML), which are then provided to the language model, alongside the natural language navigation instruction and the sensory input, as shown in Fig. 1. We make this a deliberate design choice to avoid changes to the model architecture, additional training objectives, or specialized graph encoders, thereby maintaining model-agnostic applicability and ease of integration. To the best of our knowledge, this is the first work to inject a traffic scene graph as explicit context for planning in a language-based driving model.
@@
Our experiments reveal several key insights. First, employing scene graph prompts at test-time alone improves performance over the original models. Second, training with scene graph supervision yields significant gains in driving performance, and importantly, these improvements persist even when scene graphs are omitted at test-time. This indicates that language-based planners can internalize structured relational knowledge during training and leverage it without explicit relational input at inference. Finally, we find that leaner abstractions, such as actor-only graphs with pairwise relational links, offer a strong trade-off between semantic fidelity and prompt efficiency, achieving high performance with fewer tokens.
Our main contribution is a model-agnostic approach that grounds language-based autonomous driving models in explicit relational structure through serialized traffic scene graphs, enhancing structured reasoning over spatial, regulatory, and inter-actor dependencies. We systematically evaluate our method across graph abstraction levels, serialization formats, and prompt templates with multiple state-of-the-art language-based driving models to analyze how explicit relational context influences planning behavior and performance. This reveals a novel and practical paradigm where models benefit from scene graph supervision during training but operate effectively without it at test-time, avoiding the cost and complexity of test-time scene graphs.
@@ 2 Related Work
We categorize related work into two areas: (1) language-based autonomous driving, where VLMs are used to plan from multimodal inputs and instructions, and (2) traffic scene graphs, which provide structured representations of the driving environment.
@@
Language-Based Autonomous Driving: We review language-driven approaches for trajectory planning and decision making in autonomous driving, covering both open-loop waypoint prediction on nuScenes [caesar2020nuscenes] and closed-loop evaluation using CARLA [dosovitskiy2017carla] or NavSim [dauner2024navsim]. Several works deploy VLMs as planners, mapping sensor and a navigation command directly to a sequence of future waypoints [shao2024lmdrive, winter2025bevdriver, renz2025simlingo, zhang2025vldrive, xu2025drivegpt4v2, xie2025s4driver]. LMDrive [shao2024lmdrive] presents a VLM-based end-to-end stack that fuses camera and LiDAR inputs with a natural-language navigation command. BEVDriver [winter2025bevdriver] extends this design by strengthening the vision encoder to produce richer BEV features, improving semantic and spatial grounding. Beyond direct waypoint prediction, some approaches aim to increase reasoning fidelity by incorporating chain-of-thought within a perception-prediction-planning paradigm, accepting additional latency in exchange for interpretability and more transparent decision making [kong2025vlrdriver, zeng2025futuresightdrive, hwang2025emma, tian2024drivevlm]. Efficiency-oriented systems activate the language model only when it is expected to contribute: AdaDrive [zhang2025adadrive] learns an adaptive slow-fast collaboration that decides when and to what extent the VLM should be involved.
Other approaches decouple high-level reasoning from trajectory generation. Here, the VLM proposes subgoals, constraints, or safety checks, and a non-VLM planner converts these signals into concrete trajectories [fu2025orion, chen2025solve, han2025dme, pan2024vlp]. ORION [fu2025orion], for example, predicts planning tokens with an LLM and conditions a generative planner on these tokens to produce multi-modal trajectories. Finally, several approaches transfer the world knowledge and reasoning capabilities of VLMs into training-time supervision, removing the need for an online VLM at inference while retaining its benefits [hegde2025dima, zhang2024feedback]. In parallel, retrieval-augmented (RAG) methods condition decisions on similar scenarios or past cases to improve robustness on rare or long-tail events [yuan2024ragdriver, mei2024leapad, wen2024dilu, wang2025rad, zhang2025safeauto]. Agentic formulations expose tool APIs via function calls so the model can adaptively orchestrate perception, mapping, and planning utilities [mao2024agent].
Figure 2: Scene graph construction. Each traffic scene is represented as a structured, labeled graph capturing entities (actors, objects, structure) and their relations. We define three levels of abstraction: Full (all node types and relations), Road-Level (collapsed structural detail), and Actor-Only (actors and their pairwise interactions), enabling analysis of the trade-off between relational fidelity and prompt efficiency. Dashed nodes and edges indicate optional elements that are not always present in a given scene.
@@
On the construction side, recent work builds scene graphs from video, multi-camera perception, and HD maps, progressing from rule-based extractors to learned relation predictors and lane-topology transformers, and further to multi-agent 3D urban scene graphs [malawade2022roadscene2vec, wang2024rs2g, tian2023rsg, lv2025t2sg, tian2025query, greve2024curbsg, steinke2025curbosg, zhang2025parking]. This shift reflects a move away from hand-crafted priors toward data-driven structure that adapts to diverse traffic layouts and interaction patterns. For safety reasoning, scene graphs enable spatio-temporal modeling of interactive dynamics with multi-relational GNNs and sequence models, supporting early collision prediction and pedestrian risk assessment [malawade2022roadscene2vec, malawade2022sg2vec, liu2023learning, zhou2024hktsg]. For retrieval and downstream decision support, methods based on subgraph matching, together with hybrids that use VLM-generated descriptions, identify semantically similar scenes, mine failure patterns, and return mitigation knowledge to the planner through RAG-style augmentation [tian2023rsgsearch, tian2024rsgsearchplus, tian2025query, liao2025workzones]. Graph-based trajectory prediction approaches model interactions among agents and their surrounding map context, using graph- or transformer-based encoders to capture spatial and relational dependencies for subsequent motion forecasting [mo2022multi, zhou2022hivt, jia2023hdgt, zhang2025graphad]. Distinct from prior work, we are the first to serialize the scene graph and incorporate it directly into a language-based driving model as auxiliary context for decision making and planning.
@@ 3 Method
Our method conditions a language-based model on serialized traffic scene graphs to support structured reasoning for downstream driving decisions.
@@ 3.1 Scene Graph Construction
We first construct a scene graph for each time step by detecting nodes and extracting relations, building on [malawade2022roadscene2vec].
@@
Nodes carry a unique identifier as well as a semantic class and are organized into three disjoint groups: (i) structural nodes $\mathcal{S}_{t}$, (ii) actor nodes $\mathcal{A}_{t}$, and (iii) traffic-related objects $\mathcal{O}_{t}$, so that $V_{t}=\mathcal{S}_{t}\cup\mathcal{A}_{t}\cup\mathcal{O}_{t}$. Structural nodes encode the static roadway scaffold, i.e., lanes, roads, and junctions, which provides the spatial context and connectivity for motion. Actor nodes represent road users, including the ego vehicle, which is always present, and other participants such as light vehicles (car, van, taxi, electric vehicle), heavy vehicles (truck, bus), motorcycle, bicycle, emergency vehicles, and pedestrian. These entities and their interactions define the scene dynamics that govern the ego vehicle’s driving decisions. Traffic-related objects capture regulatory context, for example traffic light, speed limit, and stop sign, which constrain admissible behavior and help anticipate required maneuvers such as stopping or yielding.
@@
The relation vocabulary $\mathcal{R}$ is partitioned into six semantic groups. Proximity relations (safety hazard, near collision, super near, very near, near, visible) quantify distance- and risk-based interactions that are key for collision avoidance. Directional relations (direct front, side front, direct rear, side rear) describe longitudinal ordering and approach geometry that affect yielding, following, or braking decisions. Lateral relations (left of, right of) encode side-by-side context that is essential for lane keeping and safe lane changes. The hierarchical relation (is in) anchors actors and objects to their structural context (e.g., a lane “is in” a road), which localizes rules and right-of-way. Topological relations (opposes, travels to, lane change) capture connectivity and permitted motions within the lane graph. Finally, the regulatory relation (controls traffic of) ties traffic devices to the structural elements they govern.
@@
We impose lightweight typing constraints on which node classes may be connected by which relations to ensure a compact and consistent graph. The hierarchical assignment builds the scene hierarchy by linking actors to lanes, lanes to roads, and, when present, roads to junctions via is in. The ego connects directly to its lane via is in, and ego-actor interactions are restricted to behaviorally relevant geometry, so only proximity, directional, and lateral relations are instantiated between actor pairs. Regulatory edges apply only to traffic lights and connect a traffic light node to the lanes it governs via controls traffic of. We do not add regulatory edges for speed limits or stop signs. Lane-to-lane relations capture the network structure and include topological relations (travels to, lane change, opposes) as well as lateral adjacency (to left of, to right of).
[…]

### v3

@@ Abstract
Vision-language models have recently emerged as promising planners for autonomous driving, where success hinges on topology-aware reasoning over spatial structure and dynamic interactions from multimodal input. However, existing models are typically trained without supervision that explicitly encodes these relational dependencies, limiting their ability to infer how agents and other traffic entities influence one another from raw sensor data. In this work, we bridge this gap with a novel model-agnostic method that conditions language-based driving models on structured relational context in the form of traffic scene graphs. We serialize scene graphs at various abstraction levels and formats, and incorporate them into models via structured prompt templates, enabling systematic analysis of when and how relational supervision is most beneficial and computationally efficient. Extensive evaluations on the LangAuto and Bench2Drive benchmarks show that scene graph conditioning yields large and persistent improvements. We observe a substantial performance increase in the Driving Score of our proposed approach versus competitive LMDrive, BEVDriver, and SimLingo baselines. These results indicate that diverse architectures can effectively internalize and ground relational priors through scene graph-conditioned training, even without requiring scene graph input at test-time. Code, fine-tuned models, and our scene graph dataset are publicly available at https://github.com/iis-esslingen/GraphPilot.
@@ 1 Introduction
Figure 1: Explicit relational grounding through scene graph conditioning. We visualize three setups: SG00-PT (the original pretrained baseline), SG00-FT (the baseline fine-tuned without scene graphs), and SG10 (conditioned on scene graphs during training but absent at test-time). The significant performance gains across LMDrive, BEVDriver, and SimLingo suggest that models internalize relational structure during training, even when scene graphs are absent at test-time.
@@
Understanding complex traffic scenes remains a central challenge in autonomous driving, especially when decision-making is conditioned on natural language instructions [4]. Modern language-based driving agents operate at the intersection of perception, language, and planning, requiring not only accurate scene understanding but also the ability to reason about spatial structure, traffic rules, and interactions among dynamic actors [49]. While recent works have shown that vision-language models (VLMs) can map image and language inputs to plausible trajectories [36, 48, 33, 58, 57, 35, 34], their performance in complex environments remains limited. A key limitation is that current models typically rely on representations that do not explicitly encode relational structure, forcing them to infer safety-critical interactions implicitly from dense features [55].
@@
Scene graphs provide a structured solution to this limitation [18, 19]. By explicitly representing a traffic scene as a structured graph of entities and their relations, they encode not only what is present in the environment, but also how entities are related and may influence one another. In scene graphs, nodes represent objects such as vehicles, pedestrians, traffic lights, and roadway elements, while edges express spatial, semantic, and regulatory relations, e.g., which vehicle is in which lane, which traffic light controls which direction, and how nearby actors might interact. This structured relational view enhances semantic understanding and enables more informed planning [3].
@@
Despite their growing use in indoor navigation and instruction-following tasks [10, 32, 52, 13, 47, 30], scene graphs have only recently gained attention in outdoor autonomous driving, mainly for tasks such as risk estimation [27, 24], and trajectory prediction [16, 60]. However, their integration into language-based planning remains unexplored. Instead, most existing systems represent the scene using dense geometric or visual features, including image tokens [33, 58, 57] or BEV grids [36, 48, 12], followed by generic attention mechanisms. As a result, critical relationships between entities must be inferred implicitly. This gap motivates our central question: Can explicit relational structure of a traffic scene help language-based driving models internalize and reason over spatial and causal dependencies in traffic scenes?
@@
To address this question, we introduce a model-agnostic approach for injecting structured semantic context into language-based autonomous driving systems. We construct traffic scene graphs at each time step and serialize them into human-readable formats (text, JSON, YAML), which are then provided to the language model along with the navigation instruction and sensory input, as shown in Fig. 1. We make this a deliberate design choice to avoid changes to the model architecture, additional training objectives, or specialized graph encoders, thereby maintaining model-agnostic applicability and ease of integration. This is the first work to inject a traffic scene graph as explicit context for planning in a language-based driving model.
@@
Our experiments reveal several key insights. First, employing scene graph prompts at test-time alone improves performance over the original models. Second, training with scene graph supervision yields significant gains in driving performance, and importantly, these improvements persist even when scene graphs are omitted at test-time. This indicates that language-based planners can internalize structured relational knowledge during training and leverage it without explicit relational input at inference. Finally, we find that leaner scene graph abstractions offer a strong trade-off between semantic fidelity and prompt efficiency, achieving high performance with fewer tokens.
Our main contribution is a model-agnostic training paradigm that grounds language-based autonomous driving models in explicit relational structure via serialized traffic scene graphs. Specifically, our contributions are threefold: (i) Model-agnostic relational grounding, as we propose the first approach to inject structured traffic scene graphs into vision-language planners without requiring architectural changes or specialized encoders. (ii) Relational internalization, where we identify a novel paradigm in which models internalize spatial and causal dependencies during training to achieve significant performance gains even when scene graphs are absent at test-time. (iii) Efficient scene representation, where we define a hierarchy of abstractions and serialization formats that identify the most token-efficient representations for planning without sacrificing driving performance. We systematically validate these contributions across multiple state-of-the-art models (LMDrive, BEVDriver, SimLingo) to demonstrate the universal benefit of explicit relational conditioning.
@@ 2 Related Work
We discuss related work regarding (1) language-based autonomous driving, where VLMs are used to plan from multimodal inputs and instructions, and (2) traffic scene graphs, which provide structured driving scene representations.
@@
Language-Based Autonomous Driving: We review language-driven methods for trajectory planning and decision making in autonomous driving, covering both open-loop waypoint prediction on nuScenes [1] and closed-loop evaluation using CARLA [6] or NavSim [5]. Several works deploy VLMs as planners, mapping sensor data and navigation commands directly to a sequence of future waypoints [36, 48, 33, 58, 51, 50]. LMDrive [36] presents a VLM-based end-to-end stack that fuses camera and LiDAR inputs with a natural-language navigation command. BEVDriver [48] extends this design by strengthening the vision encoder to produce richer BEV features, improving semantic and spatial grounding. Beyond direct waypoint prediction, some approaches aim to increase reasoning fidelity by incorporating chain-of-thought within a perception-prediction-planning paradigm, accepting additional latency in exchange for interpretability and more transparent decision making [20, 54, 15, 38]. Efficiency-oriented systems activate the language model only when it is expected to contribute: AdaDrive [57] learns an adaptive slow-fast collaboration that decides when and to what extent the VLM should be involved. Other approaches decouple high-level reasoning from trajectory generation. Here, the VLM proposes subgoals, constraints, or safety checks, and a non-VLM planner converts these signals into concrete trajectories [7, 2, 11, 31]. ORION [7], for example, predicts planning tokens with an LLM and conditions a generative planner on these tokens to produce multi-modal trajectories. Finally, several approaches transfer the world-knowledge and reasoning capabilities of VLMs into training-time supervision, thereby removing the need for an online VLM at inference while retaining its benefits [12, 56]. In parallel, retrieval-augmented (RAG) methods condition decisions on similar scenarios or past experiences to improve robustness on rare or long-tail events [53, 28, 46, 45, 55].
Scene Graphs in Autonomous Driving: Traffic scene graphs represent road users, roadway structure, and their relations as a graph. Research spans multiple directions: construction from sensory and map data, safety and risk estimation, semantic scene retrieval, and trajectory prediction. On the construction side, recent work builds scene graphs from video, multi-camera perception, and HD maps, progressing from rule-based extractors to learned relation predictors and lane-topology transformers, and further to multi-agent 3D urban scene graphs [26, 44, 39, 25, 42, 9, 37, 59]. This shift reflects a move away from hand-crafted priors toward a data-driven structure that adapts to diverse traffic layouts and interaction patterns. For safety reasoning, scene graphs enable spatio-temporal modeling of interactive dynamics with multi-relational GNNs and sequence models, supporting early collision prediction and pedestrian risk assessment [26, 27, 24, 61]. For retrieval and downstream decision support, methods based on subgraph matching, together with hybrids that use VLM-generated descriptions, identify semantically similar scenes, mine failure patterns, and return mitigation knowledge to the planner through RAG-style augmentation [40, 41, 42, 22]. Graph-based trajectory prediction approaches model interactions among agents and their surrounding map context, using graph- or transformer-based encoders to capture spatial and relational dependencies for subsequent motion forecasting [29, 62, 16, 60]. Unlike prior work, we are the first to serialize the scene graph and incorporate it directly into a language-based driving model as auxiliary context for decision-making and planning.
@@
On the construction side, recent work builds scene graphs from video, multi-camera perception, and HD maps, progressing from rule-based extractors to learned relation predictors and lane-topology transformers, and further to multi-agent 3D urban scene graphs [26, 44, 39, 25, 42, 9, 37, 59]. This shift reflects a move away from hand-crafted priors toward data-driven structure that adapts to diverse traffic layouts and interaction patterns. For safety reasoning, scene graphs enable spatio-temporal modeling of interactive dynamics with multi-relational GNNs and sequence models, supporting early collision prediction and pedestrian risk assessment [26, 27, 24, 61]. For retrieval and downstream decision support, methods based on subgraph matching, together with hybrids that use VLM-generated descriptions, identify semantically similar scenes, mine failure patterns, and return mitigation knowledge to the planner through RAG-style augmentation [40, 41, 42, 22]. Graph-based trajectory prediction approaches model interactions among agents and their surrounding map context, using graph- or transformer-based encoders to capture spatial and relational dependencies for subsequent motion forecasting [29, 62, 16, 60]. Distinct from prior work, we are the first to serialize the scene graph and incorporate it directly into a language-based driving model as auxiliary context for decision making and planning.
@@ 3 Method
Our method conditions a language-based model on serialized traffic scene graphs to enable structured reasoning for downstream driving decisions.
@@ 3.1 Scene Graph Construction
We first construct a scene graph for each time step by detecting nodes and extracting relations, building on [26].
Figure 2: Scene graph construction. Each traffic scene is represented as a structured, labeled graph capturing entities (actors, objects, structure) and their relations. We define three levels of abstraction: Full (all node types and relations), Road-Level (collapsed structural detail), and Actor-Only (actors and their pairwise interactions), enabling analysis of the trade-off between relational fidelity and prompt efficiency. Dashed nodes and edges indicate optional elements that are not always present.
@@
Nodes carry a unique identifier as well as a semantic class and are organized into three disjoint groups: (i) structural nodes $\mathcal{S}_{t}$, (ii) actor nodes $\mathcal{A}_{t}$, and (iii) traffic-related objects $\mathcal{O}_{t}$, so that $V_{t}=\mathcal{S}_{t}\cup\mathcal{A}_{t}\cup\mathcal{O}_{t}$. Structural nodes encode the static roadway scaffold, i.e., lanes, roads, and junctions, which provide the spatial context and connectivity for motion. Actor nodes represent road users, including the ego vehicle, which is always present, and other participants such as light vehicles (car, van, taxi, electric vehicle), heavy vehicles (truck, bus), motorcycle, bicycle, emergency vehicles, and pedestrian. These entities and their interactions define the scene dynamics that govern the ego vehicle’s driving decisions. Traffic-related objects capture regulatory context, for example traffic light, speed limit, and stop sign, which constrain admissible behavior and help anticipate required maneuvers such as stopping or yielding.
@@
The relation vocabulary $\mathcal{R}$ is partitioned into six semantic groups. Proximity relations (safety hazard, near collision, super near, very near, near, visible) quantify distance- and risk-based interactions that are key for collision avoidance. Directional relations (direct front, side front, direct rear, side rear) describe longitudinal ordering and approach geometry that affect yielding, following, or braking decisions. Lateral relations (left of, right of) encode side-by-side context that is essential for lane keeping and safe lane changes. The hierarchical relation (is in) anchors actors and objects within their structural context, thereby localizing rules and right-of-way. Topological relations (opposes, travels to, lane change) define lane graph connectivity, while regulatory relations (controls traffic of) link traffic devices to the structural elements they govern.
@@
We impose lightweight typing constraints on which node classes may be connected by which relations to ensure a compact and consistent graph. The hierarchical assignment builds the scene hierarchy by linking actors to lanes, lanes to roads, and, when present, roads to junctions via is in. The ego connects directly to its lane via is in, and ego-actor interactions are restricted to behaviorally relevant geometry, so only proximity, directional, and lateral relations are instantiated between actor pairs. Regulatory edges apply only to traffic lights and connect a traffic light node to the lanes it governs via controls traffic of. We do not add regulatory edges for speed limits or stop signs. Lane-to-lane relations capture the network structure and include topological and lateral relations.
Abstraction Levels. Beyond the full graph, we introduce two abstractions to identify which structural components contribute most to planning performance, while reducing prompt length and sensitivity to extraction noise, and quantifying the trade-off between efficiency and fidelity.
Road-Level collapses fine-grained lane nodes into parent roads. It aggregates lane-level relations such as actor-road membership and travels to topological edges to the road graph while removing lane-specific relations like lane change. This reduction isolates whether high-level road geometry suffices for planning compared to dense lane-level data. The second abstraction Actor-Only removes structural and object nodes altogether and retains only the ego and other actors with their pairwise spatial and directional relations. By removing hierarchical, topological, and regulatory edges, this variant restricts the scene graph context to local inter-actor geometry to assess its contribution to the planning signal.